OpenStudy (anonymous):

A Pizza store offers three different party packages of pizza. The packages and prices are listed below. If pizzas cost the same individually or in a package, what is the cost of a medium pizza? small gathering: 1 small, 1 medium, 1 large $36 4 small, 2 medium,1 large $71 huge party: 6 small, 5 medium, 4 large $86

OpenStudy (anonymous):

the middle one is for a family reunion

OpenStudy (kainui):

Are you allowed to use matrices?

OpenStudy (anonymous):

Yes, I am!

OpenStudy (blurbendy):

you can set up these equations: S + M + L = 36 4S + 2M + L = 71 6S + 5M + 4L = 86 use your preferred method to solve elimination method seems like a good bet

OpenStudy (kainui):

You should make equations out of these 3 sentences. I'll help walk you through it, don't worry about making mistakes.

OpenStudy (kainui):

Yeah, I mean whatever

OpenStudy (anonymous):

Thank's!!